OpenStudy (anonymous):

Help Please! =( Thanks in advance =D A certain substance has a heat of vaporization of 59.42 kJ/mol. At what Kelvin temperature will the vapor pressure be 7.50 times higher than it was at 347 K? -------------------------------------------------------------------- As a hint, I am told to use the Clausius-Clapeyron equation- but I'm not sure how to go about finding the answer....

OpenStudy (anonymous):

you might be having the expression for Clausius - Clapeyron equation, given, inital pressure / final pressure enthalpy of vapourisation initial temperature you just need to plug in the values in the equation. you will get the final temperature,,, that will be your answer.

OpenStudy (aaronq):

you wanna make the ratio of pressure 7.5 so you have \[\ln \frac{ P _{1} }{ P _{2} }=\frac{ H _{vap} }{ R }(\frac{ 1 }{ T _{2} }-\frac{ 1 }{ T _{1} })\] then: P1=7.5 and P2=1 \[\ln \frac{ 7.5 }{1 }=\frac{ H _{vap} }{ R }(\frac{ 1 }{ 347 K}-\frac{ 1 }{ T _{1} })\]